OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y =3/2 x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

first you need to get equation of perpendicular line. your equation is 2y-3x=8 perpendicular equation is written by switching the coefficients and sign of x and y that is 2x-3y=k Now put the x and y coordinate of the given point and find k

OpenStudy (anonymous):

Sorry equation will be 3y+2x=k

OpenStudy (anonymous):

@georgeblue22 Am i correct??